Write an equation in slope intercept form, if possible for each line
the line through (0, -3) and (4,0).

Found 2 solutions by nerdybill, castro:
Answer by nerdybill(6958)

(0, -3) and (4,0)
.
First, find the slope:
m = (y2-y1)/(x2-x1)
m = (0-(-3))/(4-0)
m = (0+3)/(4-0)
m = 3/4
.
Next using the slope and one of the two points (0,-3), plug it into the "point-slope" form of a linear equation:
y-y1 = m(x-x1)
y-(-3) = (3/4)(x-0)
y+3 = (3/4)x
y = (3/4)x - 3
The above is your equation in the "slope-intercept" form.




Answer by castro(1)